A 6-month forward contract for corn exists with a price of $1.70 per bushel. If Farmer Jayne decides to hedge her 20,000 bushels of corn with the forward contract, what is her profit or loss if spot prices are $1.65 or $1.80 when she sells her crop in 6 months? Her total costs are $33,000.

A) $1,000 gain or $1,000 loss

B) $0 gain or $3,000 gain

C) $0 loss or $3,000 loss

D) $1,000 gain or $1,000 gain

I know that the answer is (D) but do not understand the calculations. I come up with $1,000 gain when spot price is $1.65 and a $2,000 gain when spot price is $1.80.
Answer #1

As the person is a farmer, he will be worried about fall in prices hence he will short forward
Profit of a short forward of 20000 bushels=20000*(selling price-expiry spot price)

For spot price of 1.65: Profit=20000*(1.70-1.65)=1000
For spot price of 1.80: Profit=20000*(1.70-1.80)=-2000

TOTAL PROFIT FOR THE FARMER: The farmer will sell the produce at future spot price and gain/lose on the forward position. Total gain=20000*expiry spot price+gain on futures-total cost
1.65: =20000*1.65+1000-33000=1000
1.80: =20000*1.80-2000-33000=1000

Option D)